# Token-Specific Watermarking with Enhanced Detectability and Semantic Coherence for Large Language Models

### v2

@@ 5.3 Analysis of Learned $\delta$ and $\gamma$
One observation is that when the preceding token is an adjective (ADJ) or a determiner (DET), $\gamma$ and $\delta$ tend to be assigned lower values. This pattern is notable because ADJs and DETs typically precede nouns, as detailed in Appendix F. Reducing $\gamma$ leads to the selection of fewer green tokens, while simultaneously lowering $\delta$ results in applying a smaller watermark logit to these chosen tokens. This combination effectively weakens the watermark, as it biases only a few tokens with relatively lower strength, thereby encouraging the selection of the token with the highest model logit. Applying a weaker watermark to tokens following ADJs and DETs promotes the selection of the next token with the highest model logit, which is most likely to be a noun. This approach thereby enhances both semantic coherence and syntactic consistency.
Conversely, our method allocates increased values of $\gamma$ and $\delta$ following punctuation (PUNCT) tokens. Given the minimal constraints on subsequent tokens after PUNCT, as detailed in Appendix F, our model potentially exploits this flexibility to embed a stronger watermark. It does so by putting more words into the green list (via assigning higher $\gamma$ values) and enhancing the preference for these tokens through elevated $\delta$ values.
@@ 5.4 Robustness Against Attacks on Watermarks
5.5 Computational Costs
Table 1: Generation and detection speed on OPT-1.3B for generating 200 tokens, measured in seconds.
Setting Generation Detection
No watermark 3.220
KGW 3.827 0.067
Ours 3.946 0.166
We evaluate the computational time of our method for both text generation and watermark detection, and compare it to the KGW baseline and no-watermarking. Table 1 shows the results. The runtime of our method is close to the baselines.
@@ Appendix A Post-Hoc Text Detection
Here is a supplement of related works, specifically focusing on the literature for post-hoc text detection. Post-hoc text detection aims to distinguish generated texts from human-authored texts by analyzing generated texts without access to the LLMs (i.e., treating them as black boxes). These methods primarily leverage features extracted using external language models or by training models to act as detectors. Gehrmann et al. (2019) proposes the use of statistical metrics such as perplexity, entropy, and n-gram frequency to detect LLM-generated texts, as those metrics for generated texts may be different from human-written texts. Given a target text $S$, Sniffer (Li et al., 2023) uses multiple accessible language models to compute a list of perplexity for the given text $S$, and train a linear classifier based on the perplexity features. and LLMDet (Wu et al., 2023a) rely on perplexity-based features for detection. In addition, there are supervised learning methods where classifiers are trained to specifically distinguish between texts generated by humans and LLMs. Fröhling & Zubiaga (2021) and Solaiman et al. (2019) utilized SVMs and regressions based on statistical features for LLM-generated text detection, while Rodriguez et al. (2022) and Zhong et al. (2020) employed neural networks for this purpose. Other works, such as those by Solaiman et al. (2019), Ippolito et al. (2019), Guo et al. (2023), Yu et al. (2023), and the OpenAI Detector (OpenAI, 2023), involve fine-tuning a pre-trained RoBERTa (Liu et al., 2019) classifier for the detection of LLM-generated texts.
@@ Appendix D Experimental Details
The trade-off curve between true positive rate (TPR) and semantic similarity at a specified false positive rate (FPR) is plotted for our method and the baselines for comparison. In this section, we outline the procedure for plotting the trade-off curves for different methods studied in this work. For each method, we plot this curve by varying the parameters that directly influence TPR and semantic similarity among their optimal choices, as reported in their respective papers.
@@
For the SWEET method, we set $\gamma$ at 0.25 and varied $\delta$ within the range of $[1.0,3.0]$, in line with their implementation. However, we did not adopt their entropy threshold, as it is specifically tailored to code generation. Instead, we calculated a new threshold tailored to our training dataset, using the same method they described. Specifically, for each token in a training example, we measured the entropy of generating the next token using a large language model (LLM). These measurements were then used to estimate the entropy distribution. Finally, we set the threshold as the mean of this distribution. By varying these optimal parameters, we obtained corresponding TPR and semantic similarity values on the test set, which were then plotted to create the curve for SWEET.
@@ Appendix E Results on PPL and Z-Score
Figure 8 illustrates the trade-off between the average z-score and SimCSE performance, as well as between the average z-score and the oracle model perplexity (PPL), measured by OPT-2.7B. In Figure 2, we have already demonstrated that our model achieves a superior Pareto frontier in terms of TPR and SimCSE compared to the baseline method KGW. Extending this finding, Figure 8 shows that our model also surpasses the baseline methods in both average z-score and SimCSE metrics. Additionally, in terms of PPL, our method performs comparably to the baseline method. This indicates that the fluency and coherence of the watermarked text generated by our model are on par with the baselines.
@@
Appendix F Further Analysis on Influence of Preceeding Tokens on Watermark Strength
@@
In Sec. 5.3, we analyze the learned watermark logits, $\delta$, and splitting ratios, $\gamma$. Our analysis indicates that lower $\gamma$ and $\delta$ values (weaker watermark) are assigned to adjectives (ADJ) and determiners (DET), likely due to the high likelihood of a noun following. In contrast, punctuation (PUNCT) is assigned higher $\gamma$ and $\delta$ values (stronger watermark) possibly due to the absence of restrictions on the subsequent token. To further support this claim, we estimated the transition probabilities from adjectives, determiners, and punctuation to the next token’s part of speech (POS) tag using our training dataset. Notably, our observations show that a determiner is followed by a noun with a 70% probability, while an adjective precedes a noun with a 56% probability. In the case of punctuation, we observed a nearly uniform distribution over the next token’s POS tags. These observations reinforce our claims in Sec. 5.3.
Appendix G Qualitative Analysis
Prompt
$\cdots$ Typically, when the assessed value goes up dramatically for a big multi-tenant commercial building, the property owner isn’t the one on the hook for the bill — it’s the tenants.
Sylvester, who previously sat on the B.C. Assessment Board and the Downtown Vancouver Business Improvement Association, said she’s been discussing the idea with various people recently, including, just in the last week, the former chair of B.C. Assessment, and a provincial cabinet minister.
But property-tax agent Paul Sullivan said while the idea is “not without merit,” there are some
No watermark (z-score: -0.40)
issues with it, including who would be responsible.
Who is going to be responsible for paying this? Who are the landlords?” he asked. “Are they just the large commercial property owners? Are they multi-tenant buildings where the entire business is owned by the landlord?”
But Sylseter said the city could do it by creating a new low-impact category. Under her plan, which she’s presented in a report, the existing triple-net category would become a low-impact category, which would allow small businesses to still pay the higher property rates.
It will provide relief to the business community, $\cdots$
KGW (z-score: 5.77, SimCSE: 0.73)
questions about the size of the proposed subcategory.
“To my understanding it’s going to be a single income property-income tax scheme,’’ Sullivan said. “So as a result, there are going to be, quite a few, many (businesses) that probably would fall to that part instead of the (mainstream) part.
For example, Sullivan noted, under the city’s existing tax rate, a small-business operator would pay 80 per cent of the property-tax rate on an \$800,000 home, and 15 per cent on anything above.
But if Sylvester proposes a new subcategory within that tax bracket, that employee would pay less as a business has a “limited ability to grow their cash.” $\cdots$
Ours (z-score: 6.45, SimCSE: 0.85)
questions about the size of the proposed change, what it would mean to businesses that dont have large landlords, and how it would apply to all commercial-property owners.
The biggest issue is how it would apply to non-residential buildings. If you consider that the vast majority of Vancouver is for residential, what are the implications of a change that would apply to non-residential buildings only?” said Sullivan.
But Sylvester said her proposal is to put the property-tax rate on small businesses in the form of an exemption, so businesses like hers — which range from small bakeries to boutique clothing shops — would not pay triple-net.$\cdots$
@@ Appendix G Qualitative Analysis
Table 2: Qualitative result on a randomly selected test example is presented. This includes the prompt, the no-watermark text, the watermarked text from the KGW method, and the watermarked text from our method. To assess detectability, we provide the z-scores for the no-watermark text, the KGW method, and our method. The z-score for the no-watermark text is calculated under the assumption that it was generated by our method. Additionally, we include SimCSE evaluations to measure the similarity between the text generated by the KGW and our methods and the no-watermarked text.
@@
We can also qualitatively observe the impact of the token-specific watermarking employed by our method. For example, consider the sequence ‘…would apply to all commercial…’ from the third line of our method. Here, ‘apply’ follows ‘would,’ ‘to’ follows ‘apply,’ and ‘all’ follows ’to,’ with each of these tokens being marked as green. In contrast, the token succeeding ‘all,’ i.e., ’commercial,’ is marked as red. Upon calculating the splitting ratio ($\gamma$) and the watermark logit ($\delta$) for ‘would,’ ‘apply,’ ‘to,’ and ‘all,’ we find that ‘all’ is the only token associated with both a low $\gamma$ and a low $\delta$. This demonstrates that our method adapts its watermarking strength to maintain semantic coherence for significant words like ‘commercial,’ which is present in the prompt. Conversely, for tokens like ‘apply,’ ‘to,’ ‘all,’ our model enforces a stronger watermark, aiming to boost detectability.
@@
Appendix H Ablation Study: Comparasion of MOO and Weighted Sum Optimization

### v3

@@ 5.3 Analysis of Learned $\delta$ and $\gamma$
One observation is that when the preceding token is an adjective (ADJ) or a determiner (DET), $\gamma$ and $\delta$ tend to be assigned lower values. This pattern is notable because ADJs and DETs typically precede nouns, as detailed in Appendix I. Reducing $\gamma$ leads to the selection of fewer green tokens, while simultaneously lowering $\delta$ results in applying a smaller watermark logit to these chosen tokens. This combination effectively weakens the watermark, as it biases only a few tokens with relatively lower strength, thereby encouraging the selection of the token with the highest model logit. Applying a weaker watermark to tokens following ADJs and DETs promotes the selection of the next token with the highest model logit, which is most likely to be a noun. This approach thereby enhances both semantic coherence and syntactic consistency. Similarly, our method allocates increased values of $\gamma$ and $\delta$ following punctuation (PUNCT) tokens. Given the minimal constraints on subsequent tokens after PUNCT, as detailed in Appendix I, our model potentially exploits this flexibility to embed a stronger watermark. It does so by putting more words into the green list (via assigning higher $\gamma$ values) and enhancing the preference for these tokens through elevated $\delta$ values.
@@ 6 Conclusion
Impact Statement
Our watermarking method for LLMs carries significant societal implications. By improving the detection of machine-generated texts, our approach has the potential to prevent the usage of LLMs in harmful activities, such as election manipulation campaigns, the dissemination of fake news, and academic dishonesty. Furthermore, when used responsibly, watermarking algorithms contribute to the protection of intellectual property rights, benefiting companies and content creators, and reducing the risk of unauthorized use. However, we realize the ethical considerations associated with our work, as the misapplication of these tools could mistakenly label human-generated content as LLM-generated, potentially leading to accusations of plagiarism against innocent individuals.
@@ Appendix A Post-Hoc Text Detection
Here is a supplement of related works, specifically focusing on the literature for post-hoc text detection. Post-hoc text detection aims to distinguish generated texts from human-authored texts by analyzing generated texts without access to the LLMs (i.e., treating them as black boxes). These methods primarily leverage features extracted using external language models or by training models to act as detectors. Gehrmann et al. (2019) proposes the use of statistical metrics such as perplexity, entropy, and n-gram frequency to detect LLM-generated texts, as those metrics for generated texts may be different from human-written texts. Given a target text $S$, Sniffer (Li et al., 2023) uses multiple accessible language models to compute a list of perplexity for the given text $S$, and train a linear classifier based on the perplexity features. and LLMDet (Wu et al., 2023a) rely on perplexity-based features for detection. In addition, there are supervised learning methods where classifiers are trained to specifically distinguish between texts generated by humans and LLMs. Fröhling & Zubiaga (2021) and Solaiman et al. (2019) utilized SVMs and regressions based on statistical features for LLM-generated text detection, while Rodriguez et al. (2022) and Zhong et al. (2020) employed neural networks for this purpose. Other works, such as those by Solaiman et al. (2019), Ippolito et al. (2019), Guo et al. (2023b), Yu et al. (2023), and the OpenAI Detector (OpenAI, 2023), involve fine-tuning a pre-trained RoBERTa (Liu et al., 2019) classifier for the detection of LLM-generated texts.
@@ Appendix D Experimental Details
Evaluation Details on Llama-2.
For Llama-2 7B and 13B, we directly loaded the original model from Hugging Face library. For Llama-2 70B, we loaded a 4-bit quantized model  [^2] https://huggingface.co/TheBloke/Llama-2-70B-GPTQ to fit the model to a single GPU.
@@
The trade-off curve between true positive rate (TPR) and semantic similarity at a specified false positive rate (FPR) is plotted for our method and the baselines for comparison. In this section, we outline the procedure for plotting the trade-off curves for different methods studied in this work. For each method, we plot this curve by varying the parameters that directly influence TPR and semantic similarity among their appropriate choices.
@@
For the SWEET and $\text{SWEET}_{\text{NoPrompt}}$ methods, we set $\gamma$ at 0.25 and varied $\delta$ within the range of $[1.0,8.0]$. We explored higher $\delta$ values to improve detectability by favoring green token selection. However, even with increased $\delta$, a 100% TPR at 0% FPR was not achieved. Furthermore, we did not adopt their entropy threshold value directly, as it is specifically tailored to code generation. Instead, we calculated a new threshold tailored to our training dataset, using the same method they described. Specifically, for each token in a training example, we measured the entropy of generating the next token using LLM. These measurements were then used to estimate the entropy distribution. Finally, we set the threshold as the mean of this distribution. By varying these $\delta$ parameters, we obtained corresponding TPR and semantic similarity values on the test set, which were then plotted to create the curve for SWEET. Similarly, for SIR and $\text{SIR}_{\text{NoPrompt}}$, we varied $\delta$ within the range of $[0.6,2.0]$. They are more sensitive to the $\delta$ parameter. For SIR, during watermark detection, we use prompts and the generated completions to get watermark logits, and calculate z values by averaging the watermark logits of the generated tokens. For $\text{SIR}_{\text{NoPrompt}}$, we only use the generated completions to get watermark logits and calculate z values.
MultiBit is designed to encode a multi-bit message into generated text, with successful watermark detection occurring when the decoded message matches the embedded watermark. The message length is strategically chosen to maintain a predetermined FPR. We determined the FPR by calculating the percentage of human-written texts that were incorrectly decoded as containing the watermark message. We utilized the smallest message lengths that yielded FPRs no higher than 0% and 1%, specifically 9 and 7 bits, respectively. This baseline does not employ a splitting ratio $\gamma$. For the watermark logit $\delta$, we vary in the range $[0.9,10.0]$.
EXP-edit uses exponential minimum sampling which is an approach to sample from a multinomial distribution. We set the temperature to the default value of 1. Additionally, our method and other baselines employ a Top-$k$ value of 50 during generation. However, the original EXP-edit generation process did not include the Top-$k$ feature, so we modified EXP-edit to incorporate this feature for a fair comparison. We refer to this modified version as EXP-edit (Top-$k$=50). It is important to note that indistinguishable methods like EXP-edit, which are based on pseudo-random sampling during generation, cannot easily extend to other decoding strategies like greedy sampling or beam search, where no randomness is involved.
@@ Appendix E Results on PPL and Z-Score
Figure 8 illustrates the trade-off between the average z-score and SimCSE performance, as well as between the average z-score and the oracle model perplexity (PPL), measured by OPT-2.7B. In Figure 2, we have already demonstrated that our model achieves a superior Pareto frontier in terms of TPR and SimCSE. Extending this finding, Figure 8 shows that our model also surpasses KGW in both average z-score and SimCSE metrics. Additionally, in terms of PPL, our method performs comparably to the baseline methods.
Appendix F Results Based on Theoretical FPR
Figure 9: Comparison of our method with KGW using one-sided z-score thresholds estimated with theoretical FPR (denoted by the subscript ‘th’). The results under empirical FPR (denoted by the subscript ‘exp’), shown in Figure 2 of the main paper, are also included for comparison. The empirical and theoretical FPR curves are closely aligned or overlapping in most regions, indicating that they are close estimates.
We computed the theoretical FPR of our method based on the assumption used in KGW that the length of the generated text is sufficiently large. Under this assumption, the Poisson binomial distribution in our method can be approximated by a normal distribution according to the central limit theorem, enabling the use of a z-test to determine the theoretical FPR. Similarly, KGW calculates the theoretical FPR by approximating its binomial distribution with a normal distribution. Given FPR of 0.2% and 1%, the corresponding one-sided z-scores under the normal distribution are 2.88 and 2.33, respectively. The experimental results corresponding to these theoretical FPR are shown in Figure 9, where our approach demonstrated superior performance compared to KGW. It is important to note that the theoretical FPR calculations for both KGW and our method may not be entirely precise due to the approximations of distributions (Fernandez et al., 2023).
Alongside the curves corresponding to theoretical FPR, we also present curves corresponding to the empirical FPR, as shown in Figure 2 of the main paper. The empirical FPR are computed using 500 human-written texts. For KGW, we fixed $\gamma=0.25$ as the splitting ratio, and the empirical thresholds for KGW were 2.82 and 2.17 for FPR of 0.2% and 1%, respectively. The empirical thresholds of our method, for the six different initializations $(0.25,2.0),(0.25,1.75),(0.25,1.5),(0.25,1.25),(0.25,1.0),(0.1,1.0)$, are as follows. At an FPR of 0.2%, the thresholds are {2.68, 2.74, 3.11, 3.14, 2.87, 3.30}. At an FPR of 1%, the thresholds are {2.22, 2.24, 2.42, 2.31, 2.35, 2.70}. Each point in our method has its own empirical FPR because we learn different splitting ratios, which impact the detection results. We observe that both the empirical and theoretical FPR curves are very close or overlapping in most regions, indicating that they are close estimates.
Appendix G Further Discussion on SWEET
At $(\gamma,\delta)=(0.25,3.0)$, an analysis of LLM-generated texts that SWEET fails to detect at 0% FPR indicates that, on average, only 7 out of 200 tokens are high-entropy and suitable for SWEET watermarking. This limited number of watermarkable tokens diminishes SWEET’s detectability, even at high $\delta$. To illustrate this limitation, consider a hypothetical scenario: even if, in the best case scenario (which is rare), all 7 out of 7 tokens are identified as green tokens, the z-score calculated using the formula $(|s|_{G}-\gamma T)/\sqrt{T\gamma(1-\gamma)}$ is 4.58. As per SWEET’s implementation, $T$ is the total number of high-entropy tokens, and $|s|_{G}$ represents the identified green list tokens among them. However, with a higher number of watermarkable tokens, say 70, even if 50 were identified as green tokens, the z-score would increase to 8.97. Thus, the limited number of high-entropy tokens in SWEET hampers its detectability, highlighting the need for more watermarkable tokens for effective watermarking.
Appendix H Further Discussion on EXP-edit
In this section, we delve into a detailed comparison of our method with EXP-edit. EXP-edit’s generation process, which employs exponential minimum sampling, is pseudo-random. This pseudo-random sampling process enables sampling from a multinomial distribution. By using a watermark key, this pseudo-random process becomes deterministic. Detection of the watermark relies on an edit score (Kuditipudi et al., 2023), which measures the likelihood of the text being watermarked given this watermark key.
EXP-edit is considered an indistinguishable method because it does not bias the output LLM distribution towards specific tokens, instead exploiting the randomness of the sampling strategy for embedding the watermark. We examine this indistinguishable property using the perplexity (PPL) of the generated texts. We use an oracle model, OPT-2.7B, to compute PPL of generated text. PPL can estimate whether the output text is sampled from the same probability distribution, as such samples will have closer PPL values. However, it is important to note that PPL is not a reliable indicator of semantic quality. For example, the authors of Piet et al. (2023) found that PPL favored repeated texts, with one model producing a partial response followed by repeated ‘$l$’ characters, which was preferred due to its lower PPL.
Table 3: Comparison of TPR and PPL for No Watermark, EXP-edit, and Our Method on OPT-1.3B. The temperature is set to 1. Top-$k$=0 indicates sampling from the entire vocabulary. Exponential minimum sampling is an approach to sample from a multinomial distribution.
Method Decoding Strategy TPR @ 0% TPR @ 1% PPL
No Watermark Beam search (Num Beams=2) - - 1.628
Greedy decoding - - 1.761
Multinomial sampling (Top-$k$=50) - - 8.210
Multinomial sampling (Top-$k$=0) - - 13.241
EXP-edit Exponential minimum sampling (Top-$k$=50) 0.968 0.996 9.602
Exponential minimum sampling (Top-$k$=0) 0.922 0.996 16.235
Ours Beam search (Num Beams=2) 0.994 0.994 1.847
Greedy decoding 0.984 0.986 2.140
Multinomial sampling (Top-$k$=50) 1.000 1.000 12.227
The indistinguishable property of EXP-edit is evident when comparing the PPL values. Under multinomial sampling from the entire vocabulary, No watermark generation (Top-$k$=0) has a PPL of 13.241, while EXP-edit has a PPL of 16.235, as shown in Table 3. Similarly, with Top-$k$=50, the PPL of EXP-edit (9.602) is close to that of No watermark (8.210).
On the other hand, control over the decoding strategy is also necessary to meet specific goals. For instance, if the goal is to attain the lowest PPL, beam search or greedy decoding can be used. As shown, No watermark with beam search, with number of beams 2, achieves the lowest PPL of 1.628, followed by greedy decoding with a PPL of 1.761. Our method, based on KGW, biases the distribution towards a specific set of tokens for embedding the watermark and can be applied on top of any decoding strategy as a heuristic. As demonstrated, our method with beam search decoding (num beams = 2) achieves a PPL of 1.847, and with greedy decoding, a PPL of 2.140, outperforming both variants of EXP-edit. Additionally, our method achieves good TPR values at both 0% and 1% FPR. In contrast, it is not straightforward to extend EXP-edit to greedy and beam search, which do not involve randomness.
Furthermore, the authors of Piet et al. (2023) claim that indistinguishable methods are overly restrictive, and KGW-based methods offer more freedom to the user, enabling better detectability without significant loss of text quality.
@@ Appendix H Further Discussion on EXP-edit
Appendix I Further Analysis on Influence of Preceding Tokens on Watermark Strength
@@
In Sec. 5.3, we analyze the learned watermark logits, $\delta$, and splitting ratios, $\gamma$. Our analysis indicates that lower $\gamma$ and $\delta$ values (weaker watermark) are assigned to adjectives (ADJ) and determiners (DET), likely due to the high likelihood of a noun following. In contrast, punctuation (PUNCT) is assigned higher $\gamma$ and $\delta$ values (stronger watermark) possibly due to the absence of restrictions on the subsequent token. To further support this claim, we estimated the transition probabilities from adjectives, determiners, and punctuation to the next token’s part of speech (POS) tag using our training dataset. Notably, our observations show that a determiner is followed by a noun with a 0.7 probability, while an adjective is followed by a noun with a 0.56 probability. In the case of punctuation, we observed a nearly uniform distribution over the next token’s POS tags. These observations reinforce our claims in Sec. 5.3.
Appendix J Qualitative Analysis
@@
Table 4: Qualitative result on a randomly selected test example is presented. This includes the prompt, the no-watermark text, the watermarked text from the KGW method, and the watermarked text from our method. To assess detectability, we provide the z-scores for the no-watermark text, the KGW method, and our method. The z-score for the no-watermark text is calculated under the assumption that it was generated by our method. Additionally, we include SimCSE evaluations to measure the similarity between the text generated by the KGW and our methods, and the no-watermarked text.
Prompt $\cdots$ Typically, when the assessed value goes up dramatically for a big multi-tenant commercial building, the property owner isn’t the one on the hook for the bill — it’s the tenants. Sylvester, who previously sat on the B.C. Assessment Board and the Downtown Vancouver Business Improvement Association, said she’s been discussing the idea with various people recently, including, just in the last week, the former chair of B.C. Assessment, and a provincial cabinet minister. But property-tax agent Paul Sullivan said while the idea is “not without merit,” there are some No watermark (z-score: -0.40) issues with it, including who would be responsible. Who is going to be responsible for paying this? Who are the landlords?” he asked. “Are they just the large commercial property owners? Are they multi-tenant buildings where the entire business is owned by the landlord?” But Sylseter said the city could do it by creating a new low-impact category. Under her plan, which she’s presented in a report, the existing triple-net category would become a low-impact category, which would allow small businesses to still pay the higher property rates. It will provide relief to the business community, $\cdots$
KGW (z-score: 5.77, SimCSE: 0.73) questions about the size of the proposed subcategory. “To my understanding it’s going to be a single income property-income tax scheme,’’ Sullivan said. “So as a result, there are going to be, quite a few, many (businesses) that probably would fall to that part instead of the (mainstream) part. For example, Sullivan noted, under the city’s existing tax rate, a small-business operator would pay 80 per cent of the property-tax rate on an \$800,000 home, and 15 per cent on anything above. But if Sylvester proposes a new subcategory within that tax bracket, that employee would pay less as a business has a “limited ability to grow their cash.” $\cdots$ Ours (z-score: 6.45, SimCSE: 0.85) questions about the size of the proposed change, what it would mean to businesses that dont have large landlords, and how it would apply to all commercial-property owners. The biggest issue is how it would apply to non-residential buildings. If you consider that the vast majority of Vancouver is for residential, what are the implications of a change that would apply to non-residential buildings only?” said Sullivan. But Sylvester said her proposal is to put the property-tax rate on small businesses in the form of an exemption, so businesses like hers — which range from small bakeries to boutique clothing shops — would not pay triple-net.$\cdots$
@@
We can also qualitatively observe the impact of the token-specific watermarking employed by our method. For example, consider the sequence ‘…would apply to all commercial…’ from the third line of our method. Here, ‘apply’ follows ‘would,’ ‘to’ follows ‘apply,’ and ‘all’ follows ‘to,’ with each of these tokens being marked as green. In contrast, the token succeeding ‘all,’ i.e., ‘commercial,’ is marked as red. Upon calculating the splitting ratio ($\gamma$) and the watermark logit ($\delta$) for ‘would,’ ‘apply,’ ‘to,’ and ‘all,’ we find that ‘all’ is the only token associated with both a low $\gamma$ and a low $\delta$. This demonstrates that our method adapts its watermarking strength to maintain semantic coherence for significant words like ‘commercial,’ which is present in the prompt. Conversely, for tokens like ‘apply,’ ‘to,’ ‘all,’ our model enforces a stronger watermark, aiming to boost detectability.
@@
Figure 10: Comparison of multi-objective optimization and weighted sum optimization. For the weighted sum optimization, we set $\lambda_{ws}$ to $4\times 10^{-4}$ based on the averaged value of $\frac{\lambda_{moo}}{1-\lambda_{moo}}$ during the optimization.
(a) Generation length = 50 tokens.
(b) Generation length = 400 tokens.
Figure 11: Comparison of our method with KGW on short and long text generations. For short text generations, we set the generation length to 50 tokens. For long text generations, we set the generation length to 400 tokens.
Appendix K Ablation Study: Comparison of MOO and Weighted Sum Optimization
@@
Appendix L Computational Costs
Table 5: Memory utilization of OPT-1.3B for generating 200 tokens, measured in MB.
Method Generation (MB) Detection (MB)
No Watermark 3475 -
KGW 3477 503
SWEET 3477 4933
EXP-edit 3486 416
SIR 5475 2425
MultiBit 4517 1485
Ours 3731 643
We evaluate the computational time of our method for both text generation and watermark detection, and compare it to the baselines. Table 2 shows the results. We also show memory utilization in Table 5. Our method exhibits higher generation and detection speeds compared to EXP-edit, SIR, and MultiBit. The generation and detection time of our method is comparable to that of KGW, SWEET, and No Watermarking. Overall, our method achieves superior detectability and semantic coherence without a significant increase in computational cost or memory cost.
Appendix M Performance Across Varied Generation Lengths
In our main experiments, the generation length was set to 200 tokens. Here, we extend our evaluation to include both shorter and longer text generations. For short text generations, we configured the prompt length to at least 200 tokens, followed by a generation of 50 tokens by the LLM. For longer texts, the prompt was set to a minimum of 600 tokens, with the LLM generating 400 tokens. We used the C4 news-like dataset and the OPT-1.3B model for these text generations. As illustrated in Figure 11, our method is relatively better than KGW, showcasing the efficacy of our approach across various generation lengths.
Appendix N Performance on Different Datasets
(a) Essays dataset
(b) HC3 dataset
Figure 12: Comparison of our method with KGW on the Essays and HC3 datasets.
We further evaluated our learned models ($\gamma$- and $\delta$-generator networks), initially trained on the C4 dataset, on two additional datasets without further training. The Essays dataset (Schuhmann, 2023) comprises sets of instructions paired with corresponding essays. For this dataset, we used the instructions as prompts to generate essay responses. The HC3 dataset (Guo et al., 2023a) contains ChatGPT-generated text, and we use the initial 100 tokens of the ChatGPT-generated text as prompts to generate the completions. The OPT-1.3B LLM was used for the experiment, with the generation length fixed at 200 tokens. As shown in Figure 12, our method outperforms the baseline KGW method, demonstrating the robust transferability of the $\gamma$ and $\delta$ networks across different datasets.
Appendix O Modified Corruption Attack
(a) Substitution using low-$\gamma$ tokens.
(b) Substitution using high-$\gamma$ tokens.
(c) Substitution using random tokens.
Figure 13: Corruption attack where the attacker substitutes Pof tokens with low-$\gamma$, high-$\gamma$, and random tokens when the FPR is 1%. P is set to 20% and 50%.
In this evaluation, we assume the attacker knows the $\gamma$ value for each token as determined by our method, enabling them to perform a corruption attack based on this knowledge. To investigate how varying $\gamma$ values affect attack efficacy, we conducted experiments with three types of corruptions:
  • Using our trained $\gamma$ network to select 30 tokens in the vocabulary with the lowest $\gamma$ values; then randomly substituting Pof the tokens in a watermarked text with tokens uniformly selected from this low-$\gamma$ list.
  • Using our trained $\gamma$ network to select 30 tokens in the vocabulary with the highest $\gamma$ values; then randomly substituting Pof the tokens in a watermarked text with tokens uniformly selected from this high-$\gamma$ list.
  • Randomly substituting Pof the tokens in a watermarked text with tokens drawn uniformly from the vocabulary.
We tested our method using substitution ratios (P%) of 20% and 50%, and the results are shown in Figure 13. Our method consistently outperformed KGW across all attack scenarios. This underscores the robustness of our method against Corruption Attack, even when the watermarked text is corrupted using tokens selected based on $\gamma$ values derived from our trained $\gamma$ network. The strong robustness of our method mainly stems from the improved z-score, achieved through our multi-objective optimization framework that concurrently maximizes the z-score and SimCSE. As a result, even when a Corruption Attack attempts to alter the $\gamma$ scores of certain tokens, the remaining tokens maintain a sufficiently high z-score, facilitating the easy detection of the watermark.